# LLM-in-Sandbox Elicits General Agentic Intelligence

### v2

@@ Abstract
We introduce LLM-in-Sandbox, enabling LLMs to explore within a code sandbox (i.e., a virtual computer), to elicit general intelligence in non-code domains. We first demonstrate that strong LLMs, without additional training, exhibit generalization capabilities to leverage the code sandbox for non-code tasks. For example, LLMs spontaneously access external resources to acquire new knowledge, leverage the file system to handle long contexts, and execute scripts to satisfy formatting requirements. We further show that these agentic capabilities can be enhanced through LLM-in-Sandbox Reinforcement Learning (LLM-in-Sandbox-RL), which uses only non-agentic data to train models for sandbox exploration. Experiments demonstrate that LLM-in-Sandbox, in both training-free and post-trained settings, achieves robust generalization spanning mathematics, physics, chemistry, biomedicine, long-context understanding, and instruction following. Finally, we analyze LLM-in-Sandbox’s efficiency from computational and system perspectives, and open-source it as a Python package to facilitate real-world deployment.
@@ 1 Introduction
The capabilities of Large Language Models (LLMs) have been progressively unlocked through different paradigms. In-context learning showed that models could generalize to new tasks without task-specific finetuning (Brown et al., 2020). Chain-of-thought prompting then elicited reasoning by guiding models to decompose problems into steps (Wei et al., 2022). Recently, agentic frameworks empowered models to leverage diverse tools across multiple turns (Anthropic, 2025b). Following this trajectory, how can we further unlock their capabilities?
In this work, we propose LLM-in-Sandbox—enabling LLMs to explore within a code sandbox—as a promising next step along this trajectory. As shown in Figure 1, the sandbox is essentially a virtual computer with terminal capabilities, widely used by code agents such as Claude Code (Anthropic, 2025a). While it is typically used for software engineering (Jimenez et al., 2023), we argue its potential extends far beyond coding. Computers are perhaps the most versatile platform ever created, virtually any task can be accomplished through them. This versatility stems from three meta-capabilities: external resource access (e.g., the internet), file management, and code execution. We hypothesize that combining LLMs with a virtual computer may unlock their potential for general intelligence.
To validate this potential, we evaluate LLM-in-Sandbox on challenging non-code tasks. Given a task input, LLMs explore within a sandbox with basic computer capabilities across multiple turns until task completion. Remarkably, without any additional training, LLMs can spontaneously leverage the code sandbox for non-code tasks, such as installing domain-specific tools to gain new abilities, utilizing file storage to process documents beyond context limits, and executing scripts to meet formatting requirements. As a result, state-of-the-art agentic LLMs achieve substantial performance gains across mathematics, physics, chemistry, biomedicine, long-context understanding, and instruction-following.
To further advance this paradigm, we propose LLM-in-Sandbox Reinforcement Learning (LLM-in-Sandbox-RL). While strong agentic models benefit directly from LLM-in-Sandbox, weaker models often struggle: performing worse in LLM-in-Sandbox mode than in vanilla LLM mode (i.e., directly generating the output without sandbox). LLM-in-Sandbox-RL bridges this gap using only general, non-agentic data. Specifically, we use general context-based tasks (Cheng et al., 2024) where the context is pre-placed as text files in the sandbox rather than directly in the model prompt, requiring the model to explore and interact with the environment. With only outcome-based rewards (Guo et al., 2025), LLM-in-Sandbox-RL enables weaker models to excel in LLM-in-Sandbox mode, significantly outperforming their LLM mode, while also enhancing models that already possess strong agentic capabilities. Crucially, this training elicits strong generalization: leading to consistent improvements across diverse out-of-domain tasks, and even enhancing vanilla LLM mode. This suggests LLM-in-Sandbox-RL can be a general method to elicit both agentic and non-agentic intelligence across models and domains.
@@
Our contributions are summarized as follows:
  • We introduce LLM-in-Sandbox, demonstrating that strong agentic LLMs exhibit generalization capabilities to exploit code sandboxes for non-code tasks across diverse domains, without additional training (Section 2).
  • We propose LLM-in-Sandbox Reinforcement Learning, which trains LLMs to explore sandbox environments using only general non-agentic data, enhancing generalization of both agentic and non-agentic intelligence across domains (Section 3).
  • We analyze LLM-in-Sandbox’s efficiency from computational and system perspectives, and open-source it as a Python package for real-world deployment (Section 4).
2 LLM-in-Sandbox Elicits General Intelligence
The core idea of LLM-in-Sandbox is to grant LLMs access to a computer where they can freely operate to complete user-specified tasks. Specifically, computers possess three meta-capabilities that form the foundation for general task-solving:
@@ 2 LLM-in-Sandbox Elicits General Intelligence
To explore the full potential of this paradigm, our design of LLM-in-Sandbox emphasizes two principles: minimal, providing a basic code sandbox with these three capabilities, and exploratory, encouraging models to discover diverse solution strategies. In the following, we describe our sandbox environment (Section 2.1), the LLM-in-Sandbox workflow (Section 2.2), as well as experiments (Section 2.3) and analysis (Section 2.4) in general domains.
2.1 Code Sandbox
A code sandbox is a virtualized computing environment, typically an Ubuntu-based system implemented via Docker containers, that provides LLMs with terminal access and full system capabilities. Within this environment, LLMs can execute arbitrary bash commands, create and modify files, and access network resources. The containerized nature ensures isolation from the host system, enabling safe execution of model-generated code.
@@ 2.1 Code Sandbox
Within the code sandbox, we equip the model with three fundamental tools that together realize the core capabilities of a computer: (1) execute_bash for executing arbitrary terminal commands, (2) str_replace_editor for file creation, viewing, and editing, and (3) submit for indicating task completion. Specifically, execute_bash is a powerful meta-tool that enables models to install packages, run programs, and even programmatically create new tools on demand, thereby bootstrapping any additional functionality beyond the provided toolset. Detailed specifications are provided in Appendix A.
@@
3:Tools: {execute_bash, str_replace_editor, submit}
@@
2.2 LLM-in-Sandbox Workflow
Our workflow builds on the ReAct framework (Yao et al., 2022), where the model iteratively reasons and acts based on environmental feedback. As shown in Algorithm 1 (purple highlights indicate sandbox-specific components), at each turn, the model generates a tool call, receives the execution result from the sandbox, and decides the next action. This multi-turn interaction continues until the model calls submit or reaches a maximum turn limit. To accommodate diverse scenarios in general tasks, our workflow encourages free exploration and supports flexible input/output handling.
@@ 2.2 LLM-in-Sandbox Workflow
We design a system prompt that guides models to fully utilize the sandbox. First, it encourages models to leverage computational tools rather than performing calculations through natural language. Second, it emphasizes deriving answers through program execution instead of directly hardcoding results. Third, it informs models that the sandbox is a safe, isolated environment where they can freely explore diverse approaches to complete tasks. The full system prompt is provided in Appendix F.
@@
We leverage the sandbox’s file system to flexibly handle diverse input/output formats. For inputs, content can be provided not only via the model prompt but also through files. For example, for long-context understanding tasks that require reading documents, we can regard the documents as task requirements and place documents in /testbed/documents/. For outputs, the model is instructed to place the final result at a designated location (e.g., /testbed/answer.txt), containing only the final result without intermediate content. After task completion, the result is extracted from this location as the final output. This approach cleanly separates exploration from final output and naturally accommodates various data formats.
2.3 Experiments on General Domains
We conduct experiments to investigate whether sandbox access improves LLM performance on general tasks. Below we present the experimental setup and results.
Setup
We compare LLM-in-Sandbox with vanilla LLM generation (i.e., directly generating the output without sandbox) across diverse models and domains. The evaluated LLMs cover frontier proprietary, open-weight, code-specialized, and smaller general-purpose models: Claude-Sonnet-4.5-Thinking (Anthropic, 2025b), GPT-5 (Singh et al., 2025), DeepSeek-V3.2-Thinking (Liu et al., 2025), MiniMax-M2 (MiniMax, 2025), Kimi-K2-Thinking (Team et al., 2025), Qwen3-Coder-30B-A3B-Instruct (Yang et al., 2025a), and Qwen3-4B-Instruct-2507 (Yang et al., 2025a).
We test on challenging tasks in six non-code domains: Mathematics, Physics, Chemistry, Biomedicine, Long-Context Understanding, and Instruction Following. For long-context tasks, we store the input documents in the sandbox environment rather than including them in the prompt, to test the model’s ability to leverage the sandbox. Since models have internet access in the sandbox, we reframe test problems to prevent benchmark hacking and manually verify sampled trajectories to ensure valid reasoning. The detailed sandbox implementations, model configurations, and evaluation protocols are in Appendices A-C.
@@ Results
2.4 Analysis on Sandbox Utilization
To understand how models leverage the sandbox environment, we conduct a case study and quantitative analysis. Specifically, we focus on the three core capabilities of sandbox: external resource access, file management, and code execution. Code execution can serve many purposes; here we specifically track computation-oriented operations. We identify these behaviors through pattern matching on model actions: (1) external resources: network requests (e.g., curl, requests.get) and package installation (e.g., pip install); (2) file management: file I/O operations (e.g., open(), json.load) and shell commands (e.g., cat, grep); (3) computation: numerical solvers, iterative algorithms, and simulation loops. Detailed classification patterns are provided in Appendix D.
@@ 2.4.1 Case Study
We conduct a case study to illustrate how strong agentic models utilize the sandbox’s core capabilities to solve non-code tasks, where strong models refer to those that perform better in LLM-in-Sandbox mode than in LLM mode in Table2. Below, we present representative trajectory snippets and Gray text indicates our commentary.
@@
Chemistry Trajectory (External Resources Access) [Turn 1] Action: execute_bash ‘pip install rdkit-pypi -q’
@@
[Turn 8] Action: execute_bash ‘apt-get install -y default-jre’
@@
[Turn 12] Action: execute_bash ‘curl -s \
@@
[Turn 14] Action: execute_bash ‘apt-get install -y libopsin-java’
@@
[Turn 18] Action: execute_bash ‘java -jar opsin-cli.jar -o smi < input.txt’
@@
Long-Context Understanding Trajectory (File Management) [Turn 5] Action: execute_bash ‘ls -la /testbed/documents/’
@@
The-State-of-Competition.txt   29048 bytes . . . [Turn 8] Action: execute_bash ‘grep -n -i ‘infringement’ ...’
@@
252: . . . estimate the number of consumer-related infringements . . . [Turn 9] Action: execute_bash ‘sed -n ‘240,280p’ ...’
@@
[Turn 15] Action: str_replace_editor create extract_industries.py
@@
Instruction Following Trajectory (Computation) [Turn 3] Action: str_replace_editor create helper.py
@@
[Turn 12] Action: str_replace_editor create generate.py
@@
if check_constraints([s1, s2, s3]): candidates.append(…) [Turn 13] Action: execute_bash ‘python3 generate.py’
@@ 2.4.2 Quantitative Analysis
We quantify these behavioral patterns to analyze how sandbox utilization varies across different models and task domains. Using the pattern matching approach described above, we measure how frequently models invoke each of the three core capabilities during their interactions, computed as the ratio of capability invocations to total interaction turns.
@@
Figure 2 shows strong models adapt their usage patterns to task requirements. Mathematics benefits most due to high computation frequency (43.4%): models verify solutions through numerical computation. Chemistry shows the highest external resource frequency (18.4%) as models install domain-specific packages. In contrast, Biomedicine shows less stable improvements with the shortest exploration (6.5 turns), suggesting that models fail to fully leverage the sandbox environment for these tasks.
Figure 2: Sandbox behavior patterns across task domains for strong agentic models. (a)-(c): Capability usage rate, computed as capability invocations / total turns. (d): Average number of interaction turns per task.
Model Prompt Sandbox
@@
Table 3: Long-context performance, both use LLM-in-Sandbox mode but place the context differently: in prompt vs. in sandbox.
@@
In Figure 2, Long-Context tasks show high file operation frequency with minimal external resource usage, indicating that models focus on understanding local context. To further validate the benefit of file-based context handling, we compare two settings under LLM-in-Sandbox mode: placing documents directly in the prompt vs. storing them in the sandbox. As shown in Table 3, storing documents in the sandbox yields substantial gains on average, with Claude, DeepSeek, and Kimi showing the largest improvements. This suggests that LLM-in-Sandbox serves as a promising solution for handling long-context tasks by offloading extensive data to the environment. However, performance varies across models: Qwen perform worse with sandbox-based context, highlighting the need for training models to effectively explore file-based information.
@@
Table 4 compares sandbox utilization between strong and weak models. Strong models effectively leverage all three capabilities with high usage rate (6–21%), while the weak model (Qwen3-4B-Instruct) achieves far lower frequency (<<3%) despite taking nearly twice as many turns (23.7 vs. 12.6). This indicates that the weak model “wanders” in the sandbox without effective tool utilization, consuming more turns while accomplishing less.

### v3

# Computer Environments Elicit General Agentic Intelligence in LLMs

@@ Abstract
Agentic intelligence in large language models (LLMs) requires not only model intrinsic capabilities but also interactions with external environments. Equipping LLMs with computers now represents a prevailing trend. However, the computer environment’s intrinsic value has not been systematically investigated, particularly its potential to elicit general capabilities. Here we introduce LLM-in-Sandbox, which virtualizes the computer as a code sandbox with only basic functionalities, and demonstrate that this minimal setting elicits computer-based meta-capabilities for general task solving: external resource access, file management, and code execution. Without additional training, strong models achieve substantial gains (up to 15.5%) across mathematics, physics, chemistry, biomedicine, long-context understanding, and instruction following, while reducing token consumption by up to 8 times. Furthermore, we develop LLM-in-Sandbox-RL to train models exclusively on non-agentic data within the sandbox, empowering weaker models to harness the environment and internalize these interactions. Our results demonstrate that computer environments elicit general intelligence, yield efficiency gains, and can be harnessed through training, serving as a promising foundation for generalist agents.
@@ 1 Introduction
The evolution toward general agents is a driving force in artificial intelligence. Large language models (LLMs) have demonstrated increasingly versatile capabilities (Brown et al., 2020; Jaech et al., 2024; Guo et al., 2025; Anthropic, 2025b), making them a promising foundation for this pursuit. However, the intrinsic capability of the model alone does not fully constitute an agent. Intelligence is fundamentally situated; it is shaped by how a model engages with an external environment to retrieve information, take actions, and adapt to feedback (Wiener and von Neumann, 1949; Russell et al., 1995; Sutton et al., 1998). From this perspective, interacting with the environment acts as a defining architectural component: it dictates whether the model’s intrinsic power can be translated into effective task-solving across diverse domains.
In the real world, computers serve as a primary platform for such interaction, providing a universal interface for humans to execute complex intellectual tasks. This versatility makes computer environments a promising foundation for grounding LLM agents. Indeed, a wave of agentic systems such as Manus (Manus, 2025), Claude Code (Anthropic, 2025a) and OpenClaw (Openclaw Community, 2026) have begun equipping LLMs with computer environments, reflecting surging interest in this direction. However, these systems interleave basic environment interaction with high-level interfaces, pre-built capabilities, and other system-level optimizations, making the source of their performance difficult to isolate. Whether the observed gains arise primarily from the environment itself, or from the broader system design built around it, is not well understood. Moreover, whether models can learn to leverage these environments, and thereby acquire genuinely generalizable agentic intelligence, has yet to be systematically examined.
Here we introduce LLM-in-Sandbox (Figure 1) to investigate whether the computer, virtualized as a code sandbox, can elicit general agentic intelligence in LLMs. Compared to the standard inference setting where the model directly produces output without environmental interaction, which we refer to as LLM mode, LLM-in-Sandbox mode allows the model to interact with the sandbox through multi-turn tool calls. A code sandbox is an isolated virtual environment for terminal access, file management, and code execution. To disentangle the role of the environment from other confounding factors, our design follows two principles: minimal, preserving only basic computer functionalities without any domain-specific additions; and exploratory, encouraging models to freely experiment within this environment. Code sandboxes have proven effective for coding tasks (Yang et al., 2024); here we focus on non-code domains to test whether their benefits extend beyond coding. We further propose LLM-in-Sandbox-RL, a reinforcement learning method that trains models within the sandbox using only non-agentic data, to investigate whether models can leverage such environments to acquire more general capabilities through learning.
@@
While strong models benefit directly from the computer environment, weaker models often struggle, performing worse in LLM-in-Sandbox mode than in LLM mode. To examine whether training can bridge this gap, we propose LLM-in-Sandbox-RL, which applies reinforcement learning within the computer environment. The training data uses only general, non-agentic context-based tasks (Cheng et al., 2024) that do not target coding or any evaluation benchmark. Furthermore, background materials (i.e., the context) are stored as files in the environment rather than provided in the model prompt, forcing models to actively explore the environment to complete each task. Together with the minimal sandbox and the absence of any downstream-specific optimization, this design ensures that any performance gains can be more credibly attributed to the model learning to interact with the environment. Experiments show that, with only outcome-based rewards (Guo et al., 2025), even this simplified training setup generalizes across all evaluated domains, spanning both non-code tasks and software engineering; weaker models, in particular, shift from underperforming LLM mode to substantially outperforming it in LLM-in-Sandbox mode, while strong models also gain further improvements. Moreover, we find that LLM-in-Sandbox-RL also improves the model itself, even in LLM mode without environmental access, because structured decomposition and self-verification developed through environmental interaction are internalized into its text-only reasoning.
Overall, our results show the computer environment provides a systematic mechanism to elicit broader capabilities in LLMs. We observe gains in both training-free and post-trained settings across task domains and model capability levels, with post-training gains extending even to settings without environmental access. These gains also come with substantial efficiency advantages, showing that capability and efficiency can improve together. Given that these results point to computer environments as a promising foundation for next-generation LLMs, we release LLM-in-Sandbox as an open-source library, compatible with popular inference backends (e.g., vLLM (Kwon et al., 2023) and SGLang (Zheng et al., 2024)) and API-based LLMs, to support future research and real-world deployment.
2 Computer Environments Elicit General Intelligence
The core idea of LLM-in-Sandbox is to grant LLMs access to a computer environment where they can freely operate to complete user-specified tasks. Specifically, computers possess three meta-capabilities that form the foundation for general task-solving:
@@ 2 Computer Environments Elicit General Intelligence
To explore the full potential of this paradigm, our design of LLM-in-Sandbox emphasizes two principles: minimal, providing a code sandbox with only basic computer functionalities, and exploratory, encouraging models to freely experiment within this environment. In the following, we describe our sandbox environment (Section 2.1), the LLM-in-Sandbox workflow (Section 2.2), as well as experiments (Section 2.3) and analysis (Section 2.4) in general domains.
2.1 Virtualizing the Computer as a Code Sandbox
A code sandbox is a virtualized computer environment, typically an Ubuntu-based system implemented via Docker containers, that provides LLMs with terminal access and full system capabilities. Within this environment, LLMs can execute arbitrary bash commands, create and modify files, and access network resources. The containerized nature ensures isolation from the host system, enabling safe execution of model-generated code.
@@ 2.1 Virtualizing the Computer as a Code Sandbox
Within the code sandbox, we equip the model with three fundamental tools that together realize the core capabilities of a computer: (1) bash for executing arbitrary terminal commands, (2) file_editor for file creation, viewing, and editing, and (3) finish for indicating task completion. Specifically, bash is a powerful meta-tool that enables models to install packages, run programs, and even programmatically create new tools on demand, thereby bootstrapping any additional functionality beyond the provided toolset. Detailed specifications are provided in Appendix A.
2.2 LLM-in-Sandbox Workflow
Our workflow builds on the ReAct framework (Yao et al., 2023), where the model iteratively reasons and acts based on environmental feedback. As shown in Algorithm 1 (purple highlights indicate sandbox-specific components), at each turn, the model generates a tool call, receives the execution result from the sandbox, and decides the next action. This multi-turn interaction continues until the model calls finish or reaches a maximum turn limit. To accommodate diverse scenarios in general tasks, our workflow encourages free exploration and supports flexible input/output handling.
@@ 2.2 LLM-in-Sandbox Workflow
3:Tools: {bash, file_editor, finish}
@@
We design a system prompt that guides models to fully utilize the environment. First, it encourages models to leverage computational tools rather than performing calculations through natural language. Second, it emphasizes deriving answers through program execution instead of directly hardcoding results. Third, it informs models that the environment is a safe, isolated space where they can freely explore diverse approaches to complete tasks. The full system prompt is provided in Appendix F.
@@
We leverage the computer’s file system to flexibly handle diverse input/output formats. For inputs, content can be provided not only via the model prompt but also through files. For example, for long-context understanding tasks that require reading documents, we can regard the documents as task requirements and place documents in /testbed/documents/. For outputs, the model is instructed to place the final result at a designated location (e.g., /testbed/answer.txt), containing only the final result without intermediate content. After task completion, the result is extracted from this location as the final output. This approach cleanly separates exploration from final output and naturally accommodates various data formats.
@@
2.3 General Task Performance in Computer Environments
We conduct experiments to investigate whether computer environment access improves LLM performance on general tasks. Below we present the experimental setup and results.
Setup
We compare LLM-in-Sandbox with vanilla LLM generation (i.e., directly generating the output without environmental access) across diverse models and domains. The evaluated LLMs cover frontier proprietary, open-weight, code-specialized, and smaller general-purpose models: Claude-Sonnet-4.5-Thinking (Anthropic, 2025b), GPT-5 (Singh et al., 2025), DeepSeek-V3.2-Thinking (Liu et al., 2025), MiniMax-M2 (MiniMax, 2025), Kimi-K2-Thinking (Team et al., 2025), Qwen3-Coder-30B-A3B-Instruct (Yang et al., 2025a), and Qwen3-4B-Instruct-2507 (Yang et al., 2025a).
We test on challenging tasks in six domains: Mathematics, Physics, Chemistry, Biomedicine, Long-Context Understanding, and Instruction Following. For long-context tasks, we store the input documents in the sandbox rather than including them in the prompt, to test the model’s ability to leverage the environment. Since models have internet access in the sandbox, we reframe test problems to prevent benchmark hacking and manually verify sampled trajectories to ensure valid reasoning. The detailed environment implementations, model configurations, and evaluation protocols are in Appendices A-C.
@@ Results
2.4 Computer Capability Usage Across Domains and Model Strengths
To understand how models leverage the computer environment, we conduct a case study and quantitative analysis. Specifically, we find three computer-based meta-capabilities that benefit general task-solving: external resource access, file management, and code execution. Code execution can serve many purposes; here we specifically track computation-oriented operations. We identify these behaviors through pattern matching on model actions: (1) external resources: network requests (e.g., curl, requests.get) and package installation (e.g., pip install); (2) file management: file I/O operations (e.g., open(), json.load) and shell commands (e.g., cat, grep); (3) computation: numerical solvers, iterative algorithms, and simulation loops. Detailed classification patterns are provided in Appendix D.
@@ 2.4.1 Case Study
We conduct a case study to illustrate how strong agentic models utilize the computer’s core capabilities to solve non-code tasks, where strong models refer to those that perform better in LLM-in-Sandbox mode than in LLM mode in Table2. Below, we present representative trajectory snippets and Gray text indicates our commentary.
@@
Chemistry Trajectory (External Resources Access) [Turn 1] Action: bash ‘pip install rdkit-pypi -q’
@@
[Turn 8] Action: bash ‘apt-get install -y default-jre’
@@
[Turn 12] Action: bash ‘curl -s \
@@
[Turn 14] Action: bash ‘apt-get install -y libopsin-java’
@@
[Turn 18] Action: bash ‘java -jar opsin-cli.jar -o smi < input.txt’
@@
Long-Context Understanding Trajectory (File Management) [Turn 5] Action: bash ‘ls -la /testbed/documents/’
@@
The-State-of-Competition.txt   29048 bytes . . . [Turn 8] Action: bash ‘grep -n -i ‘infringement’ ...’
@@
252: . . . estimate the number of consumer-related infringements . . . [Turn 9] Action: bash ‘sed -n ‘240,280p’ ...’
@@
[Turn 15] Action: file_editor create extract_industries.py
@@
Instruction Following Trajectory (Code Execution) [Turn 3] Action: file_editor create helper.py
@@
[Turn 12] Action: file_editor create generate.py
@@
if check_constraints([s1, s2, s3]): candidates.append(…) [Turn 13] Action: bash ‘python3 generate.py’
@@ 2.4.2 Quantitative Analysis
We quantify these behavioral patterns to analyze how computer utilization varies across different models and task domains. Using the pattern matching approach described above, we measure how frequently models invoke each of the three core capabilities during their interactions, computed as the ratio of capability invocations to total interaction turns.
@@
Figure 2 shows strong models adapt their usage patterns to task requirements. Mathematics benefits most due to high computation frequency (43.4%): models verify solutions through numerical computation. Chemistry shows the highest external resource frequency (18.4%) as models install domain-specific packages. In contrast, Biomedicine shows less stable improvements with the shortest exploration (6.5 turns), suggesting that models fail to fully leverage the computer environment for these tasks.
Figure 2: Computer capability behavior patterns across task domains for strong agentic models. (a)-(c): Capability usage rate, computed as capability invocations / total turns. (d): Average number of interaction turns per task.
Model Prompt Env.
@@
Table 3: Long-context performance, both use LLM-in-Sandbox mode but place the context differently: in prompt vs. in environment.
@@
In Figure 2, Long-Context tasks show high file operation frequency with minimal external resource usage, indicating that models focus on understanding local context. To further validate the benefit of file-based context handling, we compare two settings under LLM-in-Sandbox mode: placing documents directly in the prompt vs. storing them in the sandbox environment. As shown in Table 3, storing documents in the environment yields substantial gains on average, with Claude, DeepSeek, and Kimi showing the largest improvements. This suggests that LLM-in-Sandbox serves as a promising solution for handling long-context tasks by offloading extensive data to the environment. However, performance varies across models: Qwen perform worse with environment-based context, highlighting the need for training models to effectively explore file-based information.
@@
Table 4 compares computer utilization between strong and weak models. Strong models effectively leverage all three capabilities with high usage rate (6–21%), while the weak model (Qwen3-4B-Instruct) achieves far lower frequency (<<3%) despite taking nearly twice as many turns (23.7 vs. 12.6). This indicates that the weak model “wanders” in the environment without effective tool utilization, consuming more turns while accomplishing less.